Suppose that a perfectly competitive firm faces a market price of ​$ 12 12 per​ unit, and at this price the​ upward-sloping portion of the​ firm's marginal cost curve crosses its marginal revenue curve at an output level of 1 comma 800 1,800 units. If the firm produces 1 comma 800 1,800 ​units, its average variable costs equal ​$ 7.00 7.00 per​ unit, and its average fixed costs equal ​$ 1.00 1.00 per unit.  

What is the​ firm's profit-maximizing​ (or loss-minimizing) output​ level? nothing . ​(Enter your response as a whole number long dash — include the minus sign if necessary.​)

What is the amount of its economic profits (or losses) at this output level?

What would be the firm's decision at this price/output level?
Answer #1

丕 12- 1800

ATC = AVC + AFC = 7 + 1 = 8

Profit maximising output is 1800 because P = MC at this output.

ATC is lower than price so firm is earning profit. Profit = 1800(12 - 8) = 1800 x 4 = 7200

Firm's decision is to continue production.